OpenStudy (amistre64):

spose we gots a matrix with a nonpivot column of all zeros; does the free variable represented by that comlun still have a "1" in it?

OpenStudy (amistre64):

ugh ... \begin{pmatrix}1&*&*&0&*&*&*\\0&1&*&0&*&*&*\\0&0&1&0&*&*&*\\0&0&0&0&*&*&*\\0&0&0&0&1&*&*\\0&0&0&0&0&*&*\end{pmatrix}

OpenStudy (amistre64):

column 4 would be a free variable so:\[c_4\begin{pmatrix}0\\0\\0\\0\\0\\0\\\end{pmatrix}\]or? \[c_4\begin{pmatrix}0\\0\\0\\1\\0\\0\\\end{pmatrix}\]?

OpenStudy (amistre64):

should be 7 entries, but i spose i shoulda tried a smaller mxn

OpenStudy (anonymous):

LOL I need a tutor for linear algebra so I definitely cant help you on this one :D

OpenStudy (amistre64):

im thinking since its a free variable that it should at least have a 1 in its own spot; but since it really doesnt participate in the playing feild I have to wonder

OpenStudy (amistre64):

i might be away from the keyboard (afk for the younger denizens of OS) so if i dont respond dont worry. im ok :) but then again it could just be that I died and went to kansas :/ either way, dont worry ;)